# Are Lithium-Rich Giants Binaries?: A Radial Velocity Variability Analysis of 1,400 Giants

### v2

@@ 2.1 RV variability as a binary indicator
Among the three surveys, RAVE usually has the largest RV errors, hence $w_{12}$ was selected as the $w$-parameter found comparing Gaia and GALAH measurements, specifically, $RV_{1}=RV_{\mathrm{Gaia}}$ and $RV_{2}=RV_{\mathrm{GALAH}}$. We defined our 6 classes described in Table 1 based on different intervals in the $w_{12}$-$w_{t}$ parameter space, and a degree of certainty of variability was designated for each class as high, medium, medium-low, low, and no variability. We fixed the thresholds for each type of variability based on the separation of the RVs, and whether their error bars intersected with each other independently of the epoch. The high certainty of variability class was divided into 2 classes. For the case of a strong variation in the Gaia-GALAH parameter, we defined the class 5a, whose interval takes values of $w_{12}\geq 2$ and $w_{t}\geq 5$ to fulfill a high certainty of variability. On the other hand, when $w_{12}<2$, we only established important variations and defined the class 5b for those stars with $w_{t}\geq 11$.
@@ 2.2 Calibration against close binaries and RV stds
Our classes in Table 1 were created based on the observed characteristics of a set of RV time series chosen to be representative of the variety of cases present in the training sample used in Section 2.1. However, we do not know if the stars in this training sample are confirmed binaries or single stars. Therefore, in this Section we test our classification scheme against two sets of calibrating stars: 33 confirmed spectroscopic binaries from the SB9 catalog (Pourbaix et al. 2009) and $263$ stars from the RV standard catalog of Soubiran et al. (2018). These stars all have multiple RV measurements in their respective catalog and have registered RV in Gaia DR3 and RAVE DR6 as well. We chose to assign $RV_{1}$ and $RV_{3}$ to come from Gaia and RAVE, respectively, and $RV_{2}$ is taken from the corresponding catalog. Thus, one star has more than one subset of RVs and $w$ values associated. For example, suppose that an SB9 star has $N_{r}$ RVs registered in that catalog, we can obtain $N_{r}$ subsets ($RV_{1},RV_{2,i},RV_{3}$), for $i=1,2...,N_{r}$ corresponding to each measurement in SB9, and recalling that $RV_{1}$ and $RV_{3}$ are fixed to the Gaia and RAVE measurements. The total number of subsets obtained was $188$ for the SB9 catalog and $3477$ for the RV standard stars catalog.
@@
In Figure 2 we show the distribution of $\Delta RV_{\mathrm{max}}$, the maximum of the absolute RV difference found in the SB9 and RV stds samples, and also for the sample of red giants that we analyze in Section 3. For the SB9 and the red giants, we only considered the RVs from their respective source catalogs, and the corresponding surveys (Gaia-Galah-RAVE measurements for the giants, see Section 3 for details). For the sample of RV stds, we show different combinations of the available RV measurements for that sample. The leftmost panel shows the distribution when only the RV measurements from the RV stds catalog are taken. In this first case, the distributions of the SB9 and the RV stds are well separated and do not overlap each other. This means that, with the RV precision of the RV stds source catalog, single stars can be well separated from close binaries like those in SB9 easily, just using $\Delta RV_{\mathrm{max}}$ and a few RV epochs. So if the giants that are the focus of this paper had RV measurements of that level of precision, the main question of this paper would be straightforward to solve with this method.
@@
The middle and right-hand panels of Figure 2 show what happens when the Gaia and RAVE RVs are included in the $\Delta RV_{\mathrm{max}}$ analysis of the sample of RV stds. As we add the Gaia and RAVE data the $\Delta{RV_{\mathrm{max}}}$ of the RV stds move to larger values, and the distribution progressively overlaps with the SB9 region. This means that with the RVs available for the giants, we expect the classification with the $w$ values to have an uncertainty area where RV variable and non-variable objects will suffer some degree of misclassification.
@@
We tested our classification on the control samples of RV stds and SB9 using the $w$ values for one subset per star, i.e., only using one RV from each catalog, and then did a second test using all the subsets. The top panel of Figure 3 shows how the subsets are distributed in the $w_{t}$-$w_{12}$ logarithmic space  [^1] We used $\log(w+1)$, thus the minimum value in the log space will be 0 as $\min(w)=0$ (for any $w$ defined). considering just one (arbitrary) subset per star; the purple solid line indicates the separation between the star subsets considering Class $\leq 4$ (below the line), and with high certainty of variability, i.e., Class 5a or 5b (above the line). We found that classification based on $w$ parameters indicates medium certainty of variability to no variability for about $80.2\pm 5.5\%$ of the RV standard stars ($84.0\%$ when all standard-star subsets are used), while about $84.9_{-16.0}^{+15.1}\%$ of the SB9 stars have a high certainty of variability ($94.15\%$ when all SB9 subsets are used).
@@
This means that, for about $80$ to $95\%$ of the cases, the methodology is capable of distinguishing RV-variable and RV-non-variable stars down to a level of $300$ m/s. Since the SB9 spectroscopic binaries used in this calibration have orbital periods between $\sim$10-5000 days, it can be concluded that when we apply our method to the giants with Li measurements (Section 3), we will be able to identify close companions to those giants with periods in that same regime with a success rate of $80$ to $95\%$. At the same time, there is a $20\%$ chance that the method flags as RV variable a star that is considered an RV standard down to 300 m/s, and between $5$-$15\%$ chance of misidentifying a spectroscopic binary as non-RV-variable. The former case might be produced by large RV errors masking a real RV variation facing our methodology, and the latter could be due to cases when the specific set of RV epochs entering the analysis are not sufficiently spread over the orbital period. To account for this we would have to examine one by one all our targets, which we choose not to attempt because it beats the purpose of the statistic we construct, which can be systematically applied to a large collection of stars with RV measurements.
@@
Figure 3: Top Panel: $w_{t}$-$w_{12}$ logarithmic distribution obtained for SB9 (orange) and RV-standard stars (blue) subsets with Gaia DR3, RAVE DR6 measurements, and one measurement obtained directly from SB9 or RV standard stars catalog of Soubiran et al. (2018). Bottom Panel: $w_{t}$-$w_{12}$ logarithmic distribution for the same subsets shown in the top panel, but colored according to the cluster that the K-means algorithm decided they belonged to (i.e., cyan for stars in Cluster 0 and magenta for those in Cluster 1). For both panels, we added a purple line that indicates the separation between Class $\leq 4$ and Class 5a-b according to our classification based on the $w$ values.
@@
Next, we used a clustering algorithm to check whether the RV stds and the spectroscopic binaries are clearly separated into two groups in the $w_{12}$-$w_{t}$ parameter space. We used KMeans from scikit.cluster to apply the K-means clustering to the samples in the $w$-parameter space. Since the K-Means algorithm finds which point must belong to a certain group based on their Euclidean distance to the cluster centroids and the $w$ values in the samples can be $\gtrsim 10^{2}$, we considered finding the clusters in the logarithmic space of $w_{12}$-$w_{t}$, and the coordinates of a single point $i$ are $x_{i}=(\log(w_{12_{i}}+1),\log(w_{t_{i}}+1))$. We established that two clusters must be found, and the algorithm iterates defining two cluster centroids $\mu_{j}$ ($j=0,1$) until the cluster inertia $\phi$ has been minimized (Arthur & Vassilvitskii 2007), where $\phi$ is given by:
@@
Each point $x_{i}$ belongs to the cluster $j$ whose $\mu_{j}$ is the nearest in the minimization of $\phi$. The bottom panel of Figure 3 shows the two clusters found in the logarithmic parameter space.
We find that $96.4\%$ of stars in Cluster 0 were Class $\leq 4$ ($100\%$ for all subsets) according to our threshold for the $w$-parameter space, while $100\%$ of stars in Cluster 1 were Class 5a or 5b ($93.7\%$ for all subsets). Therefore, the clustering algorithm naturally reproduces with high correspondence the separation between RV variable and non-variable stars achieved by the criteria in Table 1. This validates the use of those criteria for our purposes of identifying binaries among giants with Li measurements. It is important to note that we consider three different classification groups: Those determined by the clustering algorithm (Cluster 0 or 1), those identified through extensive RV previous studies (SB9 or RV standards), and those classified based on the variability classes outlined in Section 2.1. While the clustering algorithm shows a separation similar to our threshold for high variability (Class 5a-b vs Class $\leq 4$), not all the objects categorized in Cluster 0 and 1 necessarily correspond to SB9 and RV standards, respectively. This occurs because our variability classification inherently involves some level of error when identifying binaries and RV standards, as we show in the top panel of Figure 3.
@@ 3 The RV variability of giants and their Li content
We now turn to the central question of this work, which is to apply the methodology developed in Section 2 to determine whether there is any correlation between binarity and a high Li abundance among low-mass giants.
@@
As previously discussed, we compiled RV measurements from Gaia DR3, GALAH DR3, and RAVE DR6. The Gaia mission uses the RVS spectrograph with a resolving power of $R=\lambda/\Delta\lambda\sim 11500$ for near-infrared spectra. The Gaia DR3 median formal precision for the RV data is of the order of $\sim 1\ \mathrm{km\ s^{-1}}$ (Katz et al. 2022). GALAH uses the HERMES spectrograph at the Australian Astronomical Observatory, reaching a resolution of $R\sim 28000$ in the optical wavelength range. The RV uncertainties in the GALAH DR3 survey are typically $\lesssim 0.1\ \mathrm{km\ s^{-1}}$ (Zwitter et al. 2021). RAVE uses the 6dF multi-object spectrograph at the UK Schmidt telescope at Siding Spring in Australia, corresponding to a resolving power of $R\sim 7500$ for a wavelength range of 8410–8795 Å. The RV uncertainties in RAVE DR6 are $<2\ \mathrm{km\ s^{-1}}$ (Steinmetz et al. 2020). Gaia and GALAH report barycentric RV measurements, whereas RAVE provides heliocentric data. Since the difference between both frames is $\lesssim 15\ \mathrm{m\ s^{-1}}$ (e.g. Nidever et al. 2002; Collier Cameron et al. 2021), much less than the typical uncertainties of the RV measurements in the surveys, we ignore any correction related to the reference of the RV data.
@@
The sample of giants to be analyzed is constructed as follows. We first took all stars in Gaia DR3 with RV determinations, and crossmatched them with GALAH DR3 (so that Li measurements are available) and RAVE DR6, to produce a sample of stars with three independent RV determinations. Starting from this sample, we then selected giants following the recipe by Martell et al. (2021) (their Section 2.2), which uses the stellar parameters from GALAH. Then, we took those stars with no issues in their data, reduction, analysis, and iron abundance determination, summarized by the following flags in GALAH DR3: flag_sp == 0, flag_fe_h == 0, extinction parameter E(B-V)¡0.33 (to avoid difficulties with the analysis of spectra of stars with high extinction), and photometric quality flagged as A, B, or C in the WISE $W_{2}$ band (added in the catalog of GALAH) which is later used for classifying the evolutionary phase of the giants (see Buder et al. 2021, for further details of the parameters). Secondly, the sample is restricted to a range of effective temperature $3000$ K $\leq T_{\mathrm{eff}}\leq$ $5730$ K and surface gravity log(g)$\leq 3.2$, which enclose the expected range for red giants in the first-ascendant RGB and the He-core burning phase. Next, we identified if there were any stars belonging to the SMC or LMC by using their spatial and kinematic properties as described by Martell et al. (2021), but no members were found in our sample.
@@
We found a total of $5667$ red giants which have RV measurements with their respective errors reported by the three surveys. We also separately selected those stars that have flag_li_fe == 0 in GALAH DR3, which indicates a line detection good enough to consider a Li abundance measurement without upper limits and flagged problems (Martell et al. 2021; Buder et al. 2021). This subset contains a total of $1418$ red giants. Li abundances were obtained using the logarithmic number density defined as A(X)=$\log(N_{X}/N_{H})+12$ for an element X. From the abundances available in GALAH, i.e., [Li/Fe] and [Fe/H], we computed A(Li)=[[[[Li/Fe]]]]+[[[[Fe/H]]]]+1.05, where A(Li)=1.05 is the abundance value of the Sun from Asplund et al. (2009). The errors for [Li/Fe] and [Fe/H] are given in GALAH DR3, then the errors for the Li abundance are obtained as eA(Li)=$\sqrt{\mathrm{e[Li/Fe]}^{2}+\mathrm{e[Fe/H]}^{2}}$. We found that the mean error is eA(Li)$\sim 0.1$ and the maximum is eA(Li)$\sim 0.25$. Since the A(Li) distributions are later analyzed with a bin size larger than the maximum error (see Section 3.1), we do not apply any additional selection based on eA(Li).
@@
The range of metallicities in both samples is about $-1<$[Fe/H]$<0.5$, with an average of [Fe/H]$\sim-0.23\pm{0.24}$ for the sample of 5667 giants, and [Fe/H]$\sim-0.26\pm{0.25}$ for the subsample of 1418 giants. Only $\lesssim 4\%$ of the giants in the samples have metallicities [Fe/H]$<-1$, where 2 of them are Li-rich. Although an anticorrelation between metallicity and fraction of close-binary systems has been observed in previous studies (e.g. Moe et al. 2019), we did not consider any specific selection or analysis related to different ranges of metallicity in our samples.
@@
On the other hand, we also considered separating our sample with Li measurements by evolutionary phase. Martell et al. (2021), based on the work of Sharma et al. (2018), used a Bayesian Stellar Parameters estimator (BSTEP) to distinguish the stellar evolution phase of giants in GALAH DR3 between RGB and Red Clump (RC). Table 1 of Martell et al. (2021) presents the evolutionary phase classification of $1262$ Li-rich giants including information on their positions (RA, DEC) and Gaia DR2 identifier. We found $180$ coincidences between this sample and our giants with A(Li)¿1.5. While they do not include the specific evolutionary phase per target in their public data for the non-Li-rich giants, the outputs of the BSTEP provide a probability of the giants to be in the RC, which is flagged in the public catalog of GALAH as is_redclump_bstep. Then, based on the argument that RC stars are standard candles and using the absolute magnitude $W_{2}$ of WISE, Martell et al. (2021) provides a method to identify the evolutionary phase: RGB when is_redclump_bstep$<0.5$ and $|W_{2}+1.63|>0.80$, and RC when is_redclump_bstep0.5absent0.5$\geq$ 0.5≥ 0.5 and $|W_{2}+1.63|\leq 0.80$. From this, we were able to separate by evolutionary phase 850 non-Li-rich giants.
@@
The left-hand panel Figure of 4 shows that both samples, all the red giants and the sub-sample with Li abundances, are distributed in the range of $\log$(g) vs $\mathrm{T_{eff}}$ where the RGB and the RC are expected to have stars with a wide range of masses and metallicities. In the right-hand panel, we confirm that by showing the $1030$ giants we were able to separate by evolutionary phase (RGB or RC) along with the total sample with A(Li) measurements in the background.
@@
At the same time, we expect that our sample must contain a fraction of giants in the asymptotic giant branch (AGB) since the BSTEP method does not distinguish this phase that can overlap in some cases with the RGB and RC in the log(g)-$\mathrm{T_{eff}}$ diagram. By considering that the evolutionary timescale of low-mass stars during the RGB plus the RC is about $\gtrsim 500$ Myr, and in the AGB is $\sim 5$ Myr (Lamers & Levesque 2017), we estimate that the fraction of AGB stars in our sample is about $1\%$. Furthermore, some of these AGB giants could be Li-rich (e.g. Reddy & Lambert 2005; Casey et al. 2016; Holanda et al. 2020), Deepak & Reddy (2019) reported that less than $5\%$ of their 335 Li-rich giants could be in the early-AGB, but these candidates can be also giants descending from the RGB tip to the RC. With this upper limit, we could expect a maximum of 9 of our Li-rich giants to be AGB.
@@
Figure 4: Left panel: $\mathrm{log(g)}$-$\mathrm{T_{eff}}$ diagram for all the red giants found with radial velocities reported simultaneously in Gaia DR3, GALAH DR3, and RAVE DR6 (yellow), and for those which also have a Li abundance reported in GALAH DR3 (green). Right panel: $\mathrm{log(g)}$-$\mathrm{T_{eff}}$ diagram for RGB (pink) and RC (orange) giants in our sample of giants with Li abundance (green background).

### v3

@@ 2.1 RV variability as a binary indicator
Among the three surveys, RAVE usually has the largest RV errors, hence $w_{12}$ was selected as the $w$-parameter found comparing Gaia and GALAH measurements, specifically, $RV_{1}=RV_{\mathrm{\textit{Gaia}}}$ and $RV_{2}=RV_{\mathrm{GALAH}}$. We defined our 6 classes described in Table 1 based on different intervals in the $w_{12}$-$w_{t}$ parameter space, and a degree of certainty of variability was designated for each class as high, medium, medium-low, low, and no variability. We fixed the thresholds for each type of variability based on the separation of the RVs, and whether their error bars intersected with each other independently of the epoch. The high certainty of variability class was divided into 2 classes. For the case of a strong variation in the Gaia-GALAH parameter, we defined the class 5a, whose interval takes values of $w_{12}\geq 2$ and $w_{t}\geq 5$ to fulfill a high certainty of variability. On the other hand, when $w_{12}<2$, we only established important variations and defined the class 5b for those stars with $w_{t}\geq 11$.
@@ 2.2 Calibration against close binaries and RV stds
Our classes in Table 1 were created based on the observed characteristics of a set of RV time series chosen to be representative of the variety of cases present in the training sample used in Sect. 2.1. However, we do not know if the stars in this training sample are confirmed binaries or single stars. Therefore, in this section we test our classification scheme against two sets of calibrating stars: 33 confirmed spectroscopic binaries from the SB9 catalog (Pourbaix et al. 2009) and $263$ stars from the RV standard catalog of Soubiran et al. (2018). These stars all have multiple RV measurements in their respective catalog and have registered RV in Gaia DR3 and RAVE DR6 as well. We chose to assign $RV_{1}$ and $RV_{3}$ to come from Gaia and RAVE, respectively, and $RV_{2}$ is taken from the corresponding catalog. Thus, one star has more than one subset of RVs and $w$ values associated. For example, suppose that an SB9 star has $N_{r}$ RVs registered in that catalog, we can obtain $N_{r}$ subsets ($RV_{1},RV_{2,i},RV_{3}$), for $i=1,2...,N_{r}$ corresponding to each measurement in SB9, and recalling that $RV_{1}$ and $RV_{3}$ are fixed to the Gaia and RAVE measurements. The total number of subsets obtained was $188$ for the SB9 catalog and $3477$ for the RV standard stars catalog.
@@
In Fig. 2 we show the distribution of $\Delta RV_{\mathrm{max}}$, the maximum of the absolute RV difference found in the SB9 and RV stds samples, and also for the sample of red giants that we analyze in Sect. 3. For the SB9 and the red giants, we only considered the RVs from their respective source catalogs, and the corresponding surveys (Gaia-GALAH-RAVE measurements for the giants, see Sect. 3 for details). For the sample of RV stds, we show different combinations of the available RV measurements for that sample. The leftmost panel shows the distribution when only the RV measurements from the RV stds catalog are taken. In this first case, the distributions of the SB9 and the RV stds are well separated and do not overlap each other. This means that, with the RV precision of the RV stds source catalog, single stars can be well separated from close binaries like those in SB9 easily, just using $\Delta RV_{\mathrm{max}}$ and a few RV epochs. So if the giants that are the focus of this paper had RV measurements of that level of precision, the main question of this paper would be straightforward to solve with this method.
@@
The middle and right-hand panels of Fig. 2 show what happens when the Gaia and RAVE RVs are included in the $\Delta RV_{\mathrm{max}}$ analysis of the sample of RV stds. As we add the Gaia and RAVE data the $\Delta{RV_{\mathrm{max}}}$ of the RV stds move to larger values, and the distribution progressively overlaps with the SB9 region. This means that with the RVs available for the giants, we expect the classification with the $w$ values to have an uncertainty area where RV variable and non-variable objects will suffer some degree of misclassification.
@@
We tested our classification on the control samples of RV stds and SB9 using the $w$ values for one subset per star, that is, only using one RV from each catalog, and then did a second test using all the subsets. The top panel of Fig. 3 shows how the subsets are distributed in the $w_{t}$-$w_{12}$ logarithmic space  [^2] We used $\log(w+1)$, thus the minimum value in the log space will be 0 as $\min(w)=0$ (for any $w$ defined). considering just one (arbitrary) subset per star; the purple solid line indicates the separation between the star subsets considering class $\leq 4$ (below the line), and with high certainty of variability, that is, class 5a or 5b (above the line). We found that classification based on $w$ parameters indicates medium certainty of variability to no variability for about $80.2\pm 5.5\%$ of the RV standard stars ($84.0\%$ when all standard-star subsets are used), while about $84.9_{-16.0}^{+15.1}\%$ of the SB9 stars have a high certainty of variability ($94.15\%$ when all SB9 subsets are used).
@@
This means that, for about $80$ to $95\%$ of the cases, the methodology is capable of distinguishing RV-variable and RV-non-variable stars down to a level of $300$ m/s. Since the SB9 spectroscopic binaries used in this calibration have orbital periods between $\sim$10-5000 days, it can be concluded that when we apply our method to the giants with Li measurements (Sect. 3), we will be able to identify close companions to those giants with periods in that same regime with a success rate of $80$ to $95\%$. At the same time, there is a $20\%$ chance that the method flags as RV variable a star that is considered an RV standard down to 300 m/s, and between $5$-$15\%$ chance of misidentifying a spectroscopic binary as non-RV-variable. The former case might be produced by large RV errors masking a real RV variation facing our methodology, and the latter could be due to cases when the specific set of RV epochs entering the analysis are not sufficiently spread over the orbital period. To account for this we would have to examine one by one all our targets, which we choose not to attempt because it beats the purpose of the statistic we construct, which can be systematically applied to a large collection of stars with RV measurements.
@@
Figure 3: The top panel shows the $w_{t}$-$w_{12}$ logarithmic distribution obtained for SB9 (orange) and RV-standard star (blue) subsets with Gaia DR3, RAVE DR6 measurements, and one measurement obtained directly from SB9 or the RV standard stars catalog of Soubiran et al. (2018). The bottom panel shows the $w_{t}$-$w_{12}$ logarithmic distribution for the same subsets shown in the top panel, but colored according to the cluster that the K-means algorithm decided they belonged to (i.e., cyan for stars in Cluster 0 and magenta for those in Cluster 1). For both panels, we added a purple line that indicates the separation between class $\leq 4$ and class 5a-b according to our classification based on the $w$ values.
@@
Next, we used a clustering algorithm to check whether the RV stds and the spectroscopic binaries are clearly separated into two groups in the $w_{12}$-$w_{t}$ parameter space. We used KMeans from scikit.cluster to apply the K-means clustering to the samples in the $w$-parameter space. Since the K-Means algorithm finds which point must belong to a certain group based on their Euclidean distance to the cluster centroids and the $w$ values in the samples can be $\gtrsim 10^{2}$, we considered finding the clusters in the logarithmic space of $w_{12}$-$w_{t}$, and the coordinates of a single point $i$ are $x_{i}=(\log(w_{12_{i}}+1),\log(w_{t_{i}}+1))$. We established that two clusters must be found, and the algorithm iterates defining two cluster centroids $\mu_{j}$ ($j=0,1$) until the cluster inertia $\phi$ has been minimized (Arthur & Vassilvitskii 2007), where $\phi$ is given by
@@
Each point $x_{i}$ belongs to the cluster $j$ whose $\mu_{j}$ is the nearest in the minimization of $\phi$. The bottom panel of Fig. 3 shows the two clusters found in the logarithmic parameter space.
We find that $96.4\%$ of stars in Cluster 0 were class $\leq 4$ ($100\%$ for all subsets) according to our threshold for the $w$-parameter space, while $100\%$ of stars in Cluster 1 were class 5a or 5b ($93.7\%$ for all subsets). Therefore, the clustering algorithm naturally reproduces with high correspondence the separation between RV variable and non-variable stars achieved by the criteria in Table 1. This validates the use of those criteria for our purposes of identifying binaries among giants with Li measurements. It is important to note that we consider three different classification groups: Those determined by the clustering algorithm (Cluster 0 or 1), those identified through extensive RV previous studies (SB9 or RV standards), and those classified based on the variability classes outlined in Sect. 2.1. While the clustering algorithm shows a separation similar to our threshold for high variability (class 5a-b vs class $\leq 4$), not all the objects categorized in Cluster 0 and 1 necessarily correspond to SB9 and RV standards, respectively. This occurs because our variability classification inherently involves some level of error when identifying binaries and RV standards, as we show in the top panel of Fig. 3.
The identification of highly variable objects may also be influenced by the time interval between RV measurements. Therefore, we recognize that our thresholds for variability criteria could benefit from improvements. However, since the number of measurements per giant in our analysis is limited in the following section, imposing a stricter selection criterion on time between observations would reduce the total number of giants analyzed. Furthermore, variability is inherently constrained by the unknown set of orbital parameters that we do not attempt to resolve in this work. Nevertheless, using the two RVs with the lowest errors to define $w_{12}$ could help mitigate the issue of RV measurements being taken too closely in time.
@@ 3 The RV variability of giants and their Li content
We now turn to the central question of this work, which is to apply the methodology developed in Sect. 2 to determine whether there is any correlation between binarity and a high Li abundance among low-mass giants. As previously discussed, we compiled RV measurements from Gaia DR3, GALAH DR3, and RAVE DR6. The Gaia mission uses the RVS spectrograph with a resolving power of $R=\lambda/\Delta\lambda\sim 11500$ for near-IR spectra. The Gaia DR3 median formal precision for the RV data is on the order of $\sim 1\ \mathrm{km\ s^{-1}}$ (Katz et al. 2023). GALAH uses the HERMES spectrograph at the Australian Astronomical Observatory, reaching a resolution of $R\sim 28000$ in the optical wavelength range. The RV uncertainties in the GALAH DR3 survey are typically $\lesssim 0.1\ \mathrm{km\ s^{-1}}$ (Zwitter et al. 2021). RAVE uses the 6dF multi-object spectrograph at the UK Schmidt telescope at Siding Spring in Australia, corresponding to a resolving power of $R\sim 7500$ for a wavelength range of 8410–8795 Å. The RV uncertainties in RAVE DR6 are $<2\ \mathrm{km\ s^{-1}}$ (Steinmetz et al. 2020). Gaia and GALAH report barycentric RV measurements, whereas RAVE provides heliocentric data. Since the difference between both frames is $\lesssim 15\ \mathrm{m\ s^{-1}}$ (e.g., Nidever et al. 2002; Collier Cameron et al. 2021), much less than the typical uncertainties of the RV measurements in the surveys, we ignore any correction related to the reference of the RV data.
@@
The sample of giants to be analyzed is constructed as follows. We first took all stars in Gaia DR3 with RV determinations, and crossmatched them with GALAH DR3 (so that Li measurements are available) and RAVE DR6, to produce a sample of stars with three independent RV determinations. Starting from this sample, we then selected giants following the recipe by Martell et al. (2021) (their Sect. 2.2), which uses the stellar parameters from GALAH. Then, we took those stars with no issues in their data, reduction, analysis, and iron abundance determination, summarized by the following flags in GALAH DR3: flag_sp == 0, flag_fe_h == 0, extinction parameter E(B-V)¡0.33 (to avoid difficulties with the analysis of spectra of stars with high extinction), and photometric quality flagged as A, B, or C in the WISE $W_{2}$ band (added in the catalog of GALAH) which is later used for classifying the evolutionary phase of the giants (see Buder et al. 2021, for further details of the parameters). Secondly, the sample is restricted to a range of effective temperature $3000$ K $\leq T_{\mathrm{eff}}\leq$ $5730$ K and surface gravity log(g)$\leq 3.2$, which enclose the expected range for red giants in the first-ascendant RGB and the He-core burning phase. Next, we identified if there were any stars belonging to the SMC or LMC by using their spatial and kinematic properties as described by Martell et al. (2021), but no members were found in our sample.
@@
We found a total of $5667$ red giants which have RV measurements with their respective errors reported by the three surveys. We also separately selected those stars that have flag_li_fe == 0 in GALAH DR3, which indicates a line detection good enough to consider a Li abundance measurement without upper limits and flagged problems (Martell et al. 2021; Buder et al. 2021). This subset contains a total of $1418$ red giants. Li abundances were obtained using the logarithmic number density defined as A(X)=$\log(N_{X}/N_{H})+12$ for an element X. From the abundances available in GALAH, that is, [Li/Fe] and [Fe/H], we computed A(Li)=[[[[Li/Fe]]]]+[[[[Fe/H]]]]+1.05, where A(Li)=1.05 is the abundance value of the Sun from Asplund et al. (2009). The errors for [Li/Fe] and [Fe/H] are given in GALAH DR3, then the errors for the Li abundance are obtained as eA(Li)=$\sqrt{\mathrm{e[Li/Fe]}^{2}+\mathrm{e[Fe/H]}^{2}}$. We found that the mean error is eA(Li)$\sim 0.1$ and the maximum is eA(Li)$\sim 0.25$. Since the A(Li) distributions are later analyzed with a bin size larger than the maximum error (see Sect. 3.1), we do not apply any additional selection based on eA(Li).
@@
The range of metallicities in both samples is about $-1<$[Fe/H]$<0.5$, with an average of [Fe/H]$\sim-0.23\pm{0.24}$ for the sample of 5667 giants, and [Fe/H]$\sim-0.26\pm{0.25}$ for the subsample of 1418 giants. Only $\lesssim 4\%$ of the giants in the samples have metallicities [Fe/H]$<-1$, where 2 of them are Li-rich. Although an anticorrelation between metallicity and fraction of close-binary systems has been observed in previous studies (e.g., Moe et al. 2019), we did not consider any specific selection or analysis related to different ranges of metallicity in our samples.
@@
On the other hand, we also considered separating our sample with Li measurements by evolutionary phase. Martell et al. (2021), based on the work of Sharma et al. (2018), used a Bayesian stellar parameters estimator (BSTEP) to distinguish the stellar evolution phase of giants in GALAH DR3 between RGB and the red clump (RC). Table 1 of Martell et al. (2021) presents the evolutionary phase classification of $1262$ Li-rich giants including information on their positions (RA, DEC) and Gaia DR2 identifier. We found $180$ coincidences between this sample and our giants with A(Li)¿1.5. While they do not include the specific evolutionary phase per target in their public data for the non-Li-rich giants, the outputs of the BSTEP provide a probability of the giants to be in the RC, which is flagged in the public catalog of GALAH as is_redclump_bstep. Then, based on the argument that RC stars are standard candles and using the absolute magnitude $W_{2}$ of WISE, Martell et al. (2021) provides a method to identify the evolutionary phase: RGB when is_redclump_bstep$<0.5$ and $|W_{2}+1.63|>0.80$, and RC when is_redclump_bstep0.5absent0.5$\geq$ 0.5≥ 0.5 and $|W_{2}+1.63|\leq 0.80$. From this, we were able to separate by evolutionary phase 850 non-Li-rich giants.
@@
The left-hand panel of Fig. 4 shows that both samples, all the red giants and the subsample with Li abundances, are distributed in the range of $\log$(g) vs $\mathrm{T_{eff}}$ where the RGB and the RC are expected to have stars with a wide range of masses and metallicities. In the right-hand panel, we confirm that by showing the $1030$ giants we were able to separate by evolutionary phase (RGB or RC) along with the total sample with A(Li) measurements in the background.
@@
At the same time, we expect that our sample must contain a fraction of giants in the asymptotic giant branch (AGB) since the BSTEP method does not distinguish this phase that can overlap in some cases with the RGB and RC in the log(g)-$\mathrm{T_{eff}}$ diagram. By considering that the evolutionary timescale of low-mass stars during the RGB plus the RC is about $\gtrsim 500$ Myr, and in the AGB is $\sim 5$ Myr (Lamers & Levesque 2017), we estimate that the fraction of AGB stars in our sample is about $1\%$. Furthermore, some of these AGB giants could be Li-rich (e.g., Reddy & Lambert 2005; Casey et al. 2016; Holanda et al. 2020), Deepak & Reddy (2019) reported that less than $5\%$ of their 335 Li-rich giants could be in the early-AGB, but these candidates can be also giants descending from the RGB tip to the RC. With this upper limit, we could expect a maximum of 9 of our Li-rich giants to be AGB.
Additionally, we use the WISE photometry (Cutri & et al. 2012) to test the IR excess in our sample. For this, we consider the color $W_{1}-W_{4}>0.5$ as an indicator of IR excess. We found that $\sim 30\%$ of the giants had IR excess regardless of their evolutionary stage or Li abundance. However, Rebull et al. (2015) discussed the implications of IR excesses and Li abundances using WISE data, pointing out that Li-rich giants with reported IR excesses may be artifacts in the catalog or sources of confusion. Then, by taking only the clean WISE data for the magnitudes $W_{1}$ and $W_{4}$ (they have the cc_flags confusion flag set to $0$ and the ph_qual photometric quality flag set to A), as done by Martell et al. (2021), we found only 4 of 150 clean giants with IR excess, 2 of them in the RGB phase and 1 RC (the remaining target could not be classified by evolutionary stage). None of these clean giants are Li-rich.
@@
Figure 4: The left panel shows the $\mathrm{log(g)}$-$\mathrm{T_{eff}}$ diagram for all the red giants found with RVs reported simultaneously in Gaia DR3, GALAH DR3, and RAVE DR6 (yellow), and for those which also have a Li abundance reported in GALAH DR3 (green). The right panel shows the $\mathrm{log(g)}$-$\mathrm{T_{eff}}$ diagram for RGB (pink) and RC (orange) giants in our sample of giants with Li abundance (green background).